OpenStudy (anonymous):

what is the slope intercept equation of the line below? WILL GIVE METAL AND BECOME FAN

OpenStudy (anonymous):

@Kuoministers @UnkleRhaukus

OpenStudy (anonymous):

@Hero @Directrix

OpenStudy (anonymous):

I know the equation written out is y= -1 + 2/5

OpenStudy (unklerhaukus):

\[\boxed{y=mx+b}\] \(m\) is the slope \(b\) is the \(y\)- intercept

OpenStudy (anonymous):

yes @UnkleRhaukus so then what do I do? :)

OpenStudy (unklerhaukus):

well your missing an \(x\)

OpenStudy (anonymous):

Oh! * -1x

OpenStudy (anonymous):

y= -1x + 2/5

OpenStudy (unklerhaukus):

right, but you don't have to write the one

OpenStudy (anonymous):

Oh but then how do I write it? :)

OpenStudy (unklerhaukus):

well you could just write y=-x+2/5

OpenStudy (anonymous):

thank you!!